# On a problem of Nathanson on non-minimal additive complements

### v1

@@ Abstract.
Let $C$ and $W$ be two sets of integers. If $C+W=\mathbb{Z}$, then $C$ is called an additive complement to $W$. We further call $C$ a minimal additive complement to $W$ if no proper subset of $C$ is an additive complement to $W$. Answering a problem of Nathanson in part, we give sufficient conditions of $W$ which has no minimal additive complements. Our result also extends the prior result of Chen and Yang.
@@ 1. Introduction
It remains to deal with the case $\inf W>-\infty$ or $\sup W<+\infty$. Without loss of generality, we may assume that $\inf W>-\infty$ since otherwise we could consider $-W$ instead of $W$. Suppose now that $\inf W=w_{0}>-\infty$, then we can consider the set
@@
Hence the criterion $\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ in Theorem B($b$) cannot be easily improved to $\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ since otherwise there will be a contradiction between Theorem B($a$) and Theorem B($b$) for this $W$. Therefore, it would be meaningful to think of the following question.
@@
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})\leq 10\quad\text{and}\quad\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})\leq 8.$
@@
Therefore, we know that $W$ does not have a minimal additive complement, thanks to the criterion from Theorem 1 with $\overline{w_{i_{t}}}=2^{k}+8$ and $\overline{w_{i_{t+1}}}=2^{k+1}$.
@@
for those $W$ assumed in Theorem 1 which are consistent with the condition of Problem 2. We next show that the condition
@@
cannot be neglected. To see this, let $W=\bigcup\limits_{k=0}^{+\infty}[10^{k},2\times 10^{k}]$. It is clear that for this $W$ there exists a sequences $i_{1}<i_{2}<\cdots$ such that
@@
$A\ ({\rm{mod}}\ m)=\big{\{}i\in[0,m-1]:i\equiv a\ ({\rm{mod}}\ m),a\in A\big{\}}.$
Now, suppose that $W$ is an eventually periodic set with periodic $m$. By shifting a number, we may assume, without loss generality, that $W$ has the following structure:
@@
$\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})<+\infty.$
Following (1.1), we write
@@
Let $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ be as in (1.2). Suppose that $S$ is an integer set satisfying the condition that for any integer set $G$ with property that
$(-\infty,n_{G}]\subseteq G+S$
for some integer $n_{G}$ (probably negative), we have
$G\setminus\{g\}+S=G+S\quad(\forall g\in G).$
As an application of Theorem 2, we deduce the following corollary which looks a bit more clean than the theorem itself.
Corollary 1.
@@
It is clear that $Y^{(1)}$ in Corollary 1 is not finite. So, it can not be easily decided whether $W$ has a minimal additive complement by Theorem D and Theorem E.
Proof of Corollary 1 via Theorem 2.
Let $G$ be a set satisfying that
$(-\infty,n_{G}]\subseteq G+k\mathbb{N}$
for some integer $n_{G}$. Let $g$ be any fixed element of $G$, we prove that
$n\in G\setminus\{g\}+k\mathbb{N}\quad(\forall g\in G)$
for any $n$, from which it follows that
$(G\setminus\{g\})+k\mathbb{N}=G+k\mathbb{N}.$
$n=g+kn^{\prime}.$
$(-\infty,n_{G}]\subseteq G+k\mathbb{N},$
By Theorem 2, there does not exist a minimal complement to $W$.
This completes the proof of Corollary 1. ∎
@@ 2. Proof of Theorem 1
from the condition of our theorem. There exists an integer $s_{0}^{*}\geq s_{0}$ such that
@@
from the condition of our theorem, there exists an integer $s_{0}^{*}\geq s_{0}$ such that
@@
for all $s_{0},s\in L$ with $s\geq s_{0}^{*}\geq s_{0}$.
Now, the remaining arguments are the same to Case 1, so we just omit them.
@@ 3. Proof of Theorem 2
Proof of Theorem 2.
@@
$n=(j-(t+t^{\prime})m)+(x+(k+t+t^{\prime})m)\in C\setminus\{c_{0}\}+W,$
which is a contradiction. The proof of Case 1 is complete by the discussions above.
@@
Thus, we have $i+mt\in\mathcal{A}_{c_{1},w_{1}}.$ Note that $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ with $Y^{(0)}$ being a finite set and $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$. We have $\mathscr{W}=X_{m}+D\ ({\rm{mod}}\ m)$. So, we deduce
@@
$c_{0}\equiv c^{\prime}\ ({\rm{mod}}\ m).$
Then there exist $c^{\prime}\in\mathscr{C}$ and $g^{\prime}\in G_{c^{\prime}}$ so that $c_{0}=c^{\prime}+mg^{\prime}$ by the definition of $G_{c^{\prime}}$. Recall that $n_{0}=c_{0}+w_{0}$ with $c_{0}\ ({\rm{mod}}\ m)\in J$ and

### v2

@@ Abstract.
Let $C$ and $W$ be two sets of integers. If $C+W=\mathbb{Z}$, then $C$ is called an additive complement to $W$. We further call $C$ a minimal additive complement to $W$ if no proper subset of $C$ is an additive complement to $W$. Answering a problem of Nathanson in part, we give sufficient conditions of $W$ which has no minimal additive complements. Our result also extends a prior result of Chen and Yang.
@@ 1. Introduction
It therefore remains to deal with the case $\inf W>-\infty$ or $\sup W<+\infty$, thanks to Theorem A. Without loss of generality, we may assume that $\inf W>-\infty$ since otherwise we could consider $-W$ instead of $W$. Suppose now that $\inf W=w_{0}>-\infty$, then we can consider the set
@@
Hence the criterion $\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ in Theorem B($b$) cannot be easily improved to $\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ since otherwise there will be a contradiction between Theorem B($a$) and Theorem B($b$) for this $W$. Therefore, it would be meaningful to think of the following question asked by Yang (private communications).
@@
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})\leq 10\quad\text{and}\quad\liminf_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})\leq 1.$
@@
Therefore, we know that $W$ does not have a minimal additive complement, thanks to the criterion from Theorem 1 with $\overline{w_{i_{t}}}=2^{k}+8$ and $\overline{w_{i_{t+1}}}=2^{k+1}+8$.
@@
for those $W$ assumed in Theorem 1 which are consistent with the requirement of Problem 2. We next show that the condition
@@
cannot be neglected. To see this, let $W=\bigcup\limits_{k=0}^{+\infty}[10^{k},2\times 10^{k}]$. It is clear that for this $W$ there exists a sequences $i_{1}<i_{2}<\cdots$ with $\overline{w_{i_{t}}}=10^{t+1}-1$ such that
@@
$A\ ({\rm{mod}}\ m)=\big{\{}i\in[0,m-1]:i\equiv a\ ({\rm{mod}}\ m)~{}\text{for~{}some}~{}a\in A\big{\}}.$
Now, suppose that $W$ is an eventually periodic set with period $m$. By shifting a number, we may assume, without loss generality, that $W$ has the following structure:
@@
$\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})<+\infty$
(comparing this with Theorem B(b)). Following (1.1), we write
@@
Note that $Y^{(1)}$ in Theorem 2 is not finite. So, it can not be easily decided whether $W$ has a minimal additive complement by Theorem D and Theorem E.
The paper will be organized as follows: In section 2, we prove Theorem 1. In section 3, we first give a proposition and then deduce Theorem 2 from it.
@@ 2. Proof of Theorem 1
from the condition of our theorem. There exists an integer $s_{0}^{*}>s_{0}$ such that
@@
from the condition of our theorem, there exists an integer $s_{0}^{*}>s_{0}$ such that
@@
for all $s_{0},s\in L$ with $s\geq s_{0}^{*}>s_{0}$.
Now, the remaining arguments are the same to the last paragraph of Case 1 , so we just omit them.
@@ 3. Proof of Theorem 2
Theorem 2 will be deduced from the following proposition.
Proposition 1.
Let $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ be as in (1.2). Suppose that $S$ is an integer set satisfying the condition that for any integer set $G$ with property
$(-\infty,n_{G}]\subseteq G+S$
for some integer $n_{G}$ (probably negative), we have
$G\setminus\{g\}+S=G+S\quad(\forall g\in G).$
Proof.
@@
$n=(j-(t+t^{\prime})m)+(x+(k+t+t^{\prime})m)\in C\setminus\{c_{0}\}+W.$
Hence, $C\setminus\{c_{0}\}+W=\mathbb{Z}$ which is a contradiction with the minimal property of $C$.
@@
Thus, we have $i+mt\in\mathcal{A}_{c_{1},w_{1}}.$ Note that $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ with $Y^{(0)}$ being a finite set and $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$. We have $\mathscr{W}=X_{m}\cup D\ ({\rm{mod}}\ m)$. So, we deduce
@@
$c_{0}\equiv c^{\prime}\ ({\rm{mod}}\ m),$
where $c^{\prime}\in\mathscr{C}$. Then there exists $g^{\prime}\in G_{c^{\prime}}$ so that $c_{0}=c^{\prime}+mg^{\prime}$ by the definition of $G_{c^{\prime}}$. Recall that $n_{0}=c_{0}+w_{0}$ with $c_{0}\ ({\rm{mod}}\ m)\in J$ and
@@
We now carry out the proof of Theorem 2.
Proof of Theorem 2 via Proposition 1.
Let $G$ be a set satisfying
$(-\infty,n_{G}]\subseteq G+k\mathbb{N}$
for some integer $n_{G}$. Let $g$ be any fixed element of $G$. We prove that
$n\in G\setminus\{g\}+k\mathbb{N}$
for any $n\in G+k\mathbb{N}$, from which it follows that
$(G\setminus\{g\})+k\mathbb{N}=G+k\mathbb{N}.$
$n=g+kn^{\prime}.$
$(-\infty,n_{G}]\subseteq G+k\mathbb{N},$
By Proposition 1, there does not exist a minimal complement to $W$. ∎